# Pau S. Pujolas

### v1

@@ 6.2 Three Country Numerical Exercises
Table 4 presents trade weighted average tariff rates and counterfactual welfare changes for various model specifications. We report whether the model specification includes service sectors and input-output linkages, and whether optimal tariffs and welfare changes are reported relative to a free trade baseline. [^24] In Appendix B.3 we show that the results are similar if there are no aggregate trade imbalances. Our result suggests that for the trade war between the United States and China, the key driver of the results is their bilateral trade imbalance. The table also shows trade weighted average tariff rate for each specifications, where the weights for each sector are equal to trade values in the baseline data. Average Nash equilibrium tariff rates range from 9.30–11.83 percent for the United States and 10.71–17.04 percent for China. Tariff rates and welfare are lowest in the free trade baseline, when services and input-output linkages are included. This reflects that tariffs incur a spillover cost when placed on goods that serve as intermediate inputs, as they distort production in other sectors.
As expected, across all specifications, welfare is lower for the United States and tariffs are higher in the Nash equilibrium compared to unilateral tariffs — highlighting the offsetting impact of retaliation inherent in trade wars. Notably, except in the pre-trade baseline when services are excluded, a tariff war results in China setting higher tariffs on the United States than vice-versa. This reflects the scenario highlighted by Figure 5, in which each country has imbalances with other countries. This result is also in part due to the United States’ trade surplus in service sectors, and the inability of China to place tariffs on these sectors. It is possible allowing for non-tariff trade measures might counteract these effects. Also notably, across all specifications China loses a trade war with the United States; whereas, the United States only loses a trade war relative to a hypothetical free trade baseline, and only after accounting for service sectors and input-output linkages.
Table 4: Optimal Tariffs and Welfare Changes by Model Structure
Unilateral Tariffs Nash Tariffs
USA USA CHN
No No No $9.95$ $0.050$ $10.04$ $0.044$ $10.71$ $-0.059$
No Yes No $9.32$ $0.012$ $9.66$ $0.010$ $10.79$ $-0.033$
Yes No No $10.34$ $0.088$ $10.44$ $0.072$ $11.54$ $-0.155$
Yes Yes No $10.36$ $0.025$ $10.41$ $0.008$ $17.04$ $-0.138$
No No Yes $11.50$ $0.131$ $11.83$ $0.021$ $13.97$ $-0.107$
No Yes Yes $9.56$ $0.024$ $9.75$ $0.001$ $13.45$ $-0.051$
Yes No Yes $11.26$ $0.202$ $11.61$ $0.004$ $15.66$ $-0.247$
Yes Yes Yes $9.17$ $0.039$ $9.30$ $-0.003$ $14.74$ $-0.132$
7.1 Welfare Effects of 2018 Tariff War
Our results suggest the United States would gain from a trade war with China involving a move towards Nash Equilibrium tariffs. A natural follow-up question is whether the United States did, in fact, experience welfare gains as a result of the 2018 trade war. If tariffs were set non-optimally, the resulting quantitative impacts might differ substantially from those discussed in the previous section, in which tariffs were chosen to maximize welfare.
We evaluate the welfare impact of the trade war between the United States and China by calibrating the model to the 2014 baseline and then solving the counterfactual model with 2019 tariffs. In the TRAINS tariff data, many sectors experienced small tariff decreases between 2014 and 2019 as part of a longer term trend of lowering tariffs. To isolate the welfare impact of the trade war escalation in tariffs, our counterfactual includes only tariff increases between 2014 and 2019, leaving tariffs unchanged in other sectors. Table 5 reports the welfare changes that resulted from tariff increases that occurred during the Trade War. The results under "US Tariffs Only" report the welfare changes that result from U.S. tariff increases on imports from China, under the assumption that Chinese tariffs on U.S. exports remain at their pre-trade war levels. The results under "US & CHN Tariffs" report the welfare changes after accounting for Chinese tariff increases on imports from the U.S.. The results show that the United States experienced welfare losses across all model specifications, even prior to accounting for retaliation from China. Chinese tariff increases, in contrast, had negligible impacts on both U.S. and Chinese welfare, but did not lead to further welfare losses for China.
Table 5: Welfare Changes ($\Delta W$%) from US-China Trade War,
by Model Structure (2019 Tariff Rate Increases)
US Tariffs Only US & CHN Tariffs
IO Services USA CHN USA CHN
No No $-0.008$ $-0.003$ $-0.009$ $-0.003$
No Yes $-0.002$ $-0.002$ $-0.002$ $-0.002$
Yes No $-0.013$ $-0.009$ $-0.014$ $-0.009$
Yes Yes $-0.003$ $-0.009$ $-0.004$ $-0.008$
These results, particularly when contrasted with the results in Table 2, suggest that neither country set tariffs to maximize welfare. The correlation between the United States tariff rate increases during the trade war and the changes implied by Nash equilibrium tariffs reported in Table 2 are negative and equal to $-0.09$. The correlation for China’s tariff rate increases is somewhat higher at $0.29$. If tariff rate increases during the trade war were set primarily for political reasons, this suggests that the United States was willing to incur economic losses for political gains, whereas China’s response was closer to economically neutral. Our results on economic neutrality are consistent with those from Fetzer and Schwarz (2021) who find that China’s retaliatory tariffs were politically targeted, rather than set to mitigate the welfare losses from the United States’ tariffs. In contrast, they find that the European Union placed a larger emphasis on mitigating welfare losses from U.S. tariff hikes; proxied for by whether retaliatory tariffs were placed on products for which the United States had a revealed comparative advantage. These results are also consistent with Kim and Margalit (2021), who argue China’s response aimed at maximizing electoral damage to Trump.
Figure 8: Welfare Implications of Unilateral Tariff Rate Reduction by China
7.2 Trade Wars with Other Trading Partners
Table 6: Optimal Tariffs and Welfare Changes by Trading Partner
Unilateral Tariffs Nash Tariffs
USA USA Partner
AUS $10.63$ $0.003$ $10.77$ $-0.013$ $6.51$ $-0.022$
BRA $10.59$ $0.013$ $10.79$ $-0.002$ $12.55$ $-0.157$
CAN $15.79$ $0.296$ $16.74$ $0.028$ $9.55$ $-3.928$
CHE $11.31$ $0.006$ $11.31$ $0.004$ $7.94$ $-0.443$
CHN $12.88$ $0.148$ $13.09$ $0.048$ $18.13$ $-0.307$
EUC $12.27$ $0.176$ $12.75$ $-0.077$ $14.38$ $-0.319$
GBR $10.95$ $0.015$ $10.97$ $0.013$ $3.85$ $-0.183$
IDN $11.32$ $0.009$ $11.32$ $0.006$ $10.62$ $-0.180$
IND $11.38$ $0.017$ $11.38$ $0.017$ $11.42$ $-0.153$
JPN $12.05$ $0.046$ $12.21$ $0.003$ $8.92$ $-0.303$
KOR $11.59$ $0.036$ $11.82$ $-0.006$ $14.04$ $-0.625$
MEX $15.13$ $0.208$ $16.61$ $-0.038$ $12.20$ $-3.428$
NOR $10.55$ $0.002$ $10.61$ $-0.000$ $7.64$ $-0.151$
RUS $10.88$ $0.008$ $10.88$ $0.008$ $7.88$ $-0.092$
TUR $10.87$ $0.004$ $10.91$ $0.000$ $9.00$ $-0.131$
TWN $11.62$ $0.011$ $11.74$ $-0.002$ $11.29$ $-0.728$
Table 7: Optimal Tariffs and Welfare Changes by Trading Partner
Unilateral Tariffs Nash Tariffs
CHN CHN Partner
AUS $13.66$ $0.026$ $13.80$ $0.009$ $6.99$ $-0.822$
BRA $13.97$ $0.005$ $13.94$ $0.006$ $13.21$ $-0.140$
CAN $14.71$ $0.003$ $14.85$ $-0.017$ $6.15$ $-0.063$
CHE $13.26$ $0.001$ $13.26$ $-0.002$ $7.83$ $-0.152$
EUC $13.64$ $0.102$ $13.99$ $-0.130$ $11.33$ $-0.117$
GBR $12.97$ $0.001$ $13.11$ $-0.017$ $6.03$ $-0.027$
IDN $12.69$ $0.009$ $12.59$ $-0.015$ $10.80$ $-0.098$
IND $12.50$ $0.001$ $12.62$ $-0.009$ $13.05$ $-0.026$
JPN $13.68$ $0.009$ $14.21$ $-0.122$ $11.55$ $-0.023$
KOR $15.14$ $0.025$ $15.60$ $-0.028$ $13.30$ $-1.033$
MEX $13.72$ $0.002$ $13.78$ $-0.008$ $8.37$ $-0.054$
NOR $13.32$ $0.001$ $13.35$ $-0.001$ $6.17$ $-0.075$
RUS $10.95$ $0.006$ $11.83$ $-0.019$ $8.62$ $-0.095$
TUR $11.54$ $0.001$ $11.69$ $-0.011$ $8.87$ $0.024$
TWN $17.60$ $0.032$ $17.87$ $0.003$ $11.86$ $-3.082$
USA $17.34$ $0.019$ $18.10$ $-0.306$ $13.09$ $0.048$
8 Concluding remarks
Our findings, derived from a multi-sector, multi-region applied general equilibrium model, reveal that countries with significant trade deficits have an enhanced ability to gain economically from trade wars. Quantitatively, however, we find these gains to be minimal when applied to a trade war between the two countries with the largest bilateral trade imbalance, the United States and China. Our analysis also highlights the importance of including input-output linkages and service sectors in models of trade policy, which reduce the potential gains of a trade war for the United States.
We believe that our paper highlights the utility of global optimization heuristics for solving computationally complex problems. Economists are likely to find these tools useful for a range of problems, particularly for those where solutions are difficult to find, but less costly to verify. Future work might adapt our methodology for computing optimal unilateral and Nash equilibrium tariffs to study alternative trade policy scenarios. Only minimal adjustments would be needed to consider scenarios in which a country seeks to maximize economic damage dealt to a partner country while minimizing the impact on their own economy. This might be done, for instance, with the goal of deterring potential instigators of future trade wars. In this vein, our framework could be extended to analyze how trade imbalances affect the formation and stability of multilateral trade agreements.
Our findings suggest that global imbalances might make it more difficult to maintain international cooperation and interfere with multinational efforts to decrease tariffs. Large deficits make repudiation of tariff agreements more economically attractive to member states. Alternatively, trade deficits may present a credible threat of tariff escalations, and thereby encourage tariff reductions to avoid future trade wars. Regardless, it seems prudent for policymakers to consider the impact of global imbalances, and factors that might exacerbate them, when designing and negotiating multilateral agreements moving forward.
“The problem of external imbalances which are yet to fall to a tolerable level remains as a factor behind the growing pressure for protectionism … The problem will not only exacerbate bilateral trade and economic friction but will also imperil the multinational free trading system … and the development of the world economy itself." Ministry of Foreign Affairs of Japan, Diplomatic Blue Book, 1991.
Appendix A Appendix: Description of Genetic Algorithm
Our genetic algorithm is outlined in Algorithm A.1. In both the algorithm and this discussion, we set country indices to correspond to the problem of the United States setting tariffs on imports from China to maximize U.S. Welfare, taking as given China’s tariffs set on imports from the United States. We do this to aid in readability. The algorithm follows identically after updating countries indices for the reverse problem (optimal tariffs for China) and for applications involving trade wars between two arbitrary countries.
Algorithm A.1 Genetic Algorithm for Computing Best Response Tariff Rates
2:  Set bounds for elasticities $\boldsymbol{\tau}_{USA}\in[\underline{\tau},\overline{\tau}]=[1.00,5.00]$ (i.e., max rate of 400 percent)
4:  Estimation of Best Response Tariff Rates
10:        Include top $P_{elite}=16$ ranked candidates as elite children
11:     end Fitness Scaling
12:     Roulette Selection: Parents are randomly selected for crossover and mutation with probabililty proportional to $F^{\{p,g\}}$
14:        Randomly select two parents, $p_{1}$ and $p_{2}$, for each child
15:        Create child $c$ from its parents according to Laplace crossover (32)
16:     end Crossover
18:        Select one parent, $p$, for each child
20:     end Mutation
21:     Increment $g$ by one and set ${\{p,g\}}={\{c,g-1\}}$.
22:     Evaluate the fitness of each candidate, $W_{USA}^{\{p,g\}}$
26:        Increment $G_{stall}$ by 1
27:     else
28:        Reset $G_{stall}=0$
29:     end if
30:     Increment $G_{total}$ by 1
31:  end while
We create $P_{X}=116$ children from crossover using Laplace crossover (Chuang et al., 2015). Laplace crossover involves selecting two parents for each child, $p_{1}$ and $p_{2}$. We implement the selection of parents according to inverse rank fitness scaling and roulette selection. Inverse rank fitness scaling involves computing $F^{\{p,g\}}\equiv 1/\sqrt{r^{\{p,g\}}}$, where $r^{\{p,g\}}$ is the rank of candidate $p$ in generation $g$, and is used to encourage exploration by ensuring selection is not restricted to only the highest fitness candidates. Roulette selection means that each candidate will be selected with probability in proportion to its scaled fitness score. For a given child with multiple parents, each parent is selected without replacement. Across children, however, parents are selected with mutation. We use this same selection procedure for selecting parent candidates for mutation, which we discuss later in this section.
${\tau_{USA,CHN}^{s}}^{\{c,g\}}$ $={\tau_{USA,CHN}^{s}}^{\{p_{1},g\}}$ (32)
$f(x|a,b)=\frac{1}{2b}e^{-{\left|x-a\right|}/b}.$
The direction of mutation (increase or decrease) is likewise determined probabilistically, based on the scaled distance of each parameter from constraint bounds. Parameters closer to their bounds more likely to move away from them, up to a maximum rate of 0.75 of the time. For this, we defined a component-wise indicator function $\mathbb{1}_{dir}^{{\{p,g\}}}$ that is equal to 1 if a random standard uniform draw, $u_{dir}$, is less than the scaled distance of the component from the constraint bounds, i.e.,
We further accelerate convergence of the genetic algorithm by using a hybrid approach that combines the genetic algorithm with a local optimization subroutine. To implement a hybrid genetic algorithm, we follow the same steps as in the genetic algorithm described by Algorithm A.1, however, we lower the stall limit to 10 and the total limit to 200. We then apply a local optimization subroutine to each of the top 16 elite candidates from the final generation of the genetic algorithm. We then re-initialize the genetic algorithm, seeding it with the same top 16 elites and with each unique solution of the local optimization subroutine that yielded improved welfare. We continue this process until we reach convergence as determined by $\max{(\lVert\Delta\boldsymbol{\tau}_{i}^{*}\rVert,\lVert\Delta W_{i}^{*}\rVert)}\geq\textit{tol}$, which more closely matches the termination criteria of Algorithm 1. This hybrid approach allows us to exploit the genetic algorithm’s ability to explore the solution space and avoid local optima, while also exploiting the local optimization subroutine’s ability to refine solutions. Even without the hybrid approach, choosing smaller stall limits and re-initializing the algorithm multiple times can yield improved performance. As before, we do not emphasize the importance of any particular local optimization heuristic. We implement a interior trust region local optimization algorithm with bounds based on Coleman and Li (1996), and find it speeds convergence and accuracy. Attempting to bypass the genetic algorithm altogether, however, and relying solely on these local optimization heuristics can be expected to lead to poor performance for models with more than a small number of sectors.
Appendix B Appendix: Robustness and Extensions
@@ B.1 Uniform Tariff Rates Across Sectors
Our methodology can be adapted to evaluate the implications of tariff rate regimes that restrict tariff rates to be uniform across sectors. To do so, we equalize tariff rates across sectors by setting $\tau_{USA,CHN}^{s}=\tau_{USA,CHN}$, and apply our prior methodology to compute optimal unilateral and Nash equilibrium tariffs, given this restriction. Table B.1 reports optimal unilateral and Nash equilibrium tariffs when tariffs are restricted to be uniform for both the United States and China. The results are similar to Table 2, including the conclusion that the United States gains from a trade war relative to the pre-trade war baseline, but experiences small welfare losses relative to a free-trade baseline.
@@
Table B.2 reports optimal unilateral and Nash equilibrium tariffs for the United States with and without the restriction that China is forced to enact uniform tariffs itself. These results show that — for the specification with input-output linkages, but without services — uniformity achieves a large majority of the potential welfare gains from optimal tariff setting compared to Table 4, where we allow tariffs to vary across sectors. Notably, the welfare implications are nearly identical regardless of whether we restrict China to implementing uniform tariffs or not. Although China does have higher welfare, and the United States has lower welfare, under the Nash equilibrium when China is not restricted to uniform tariffs, the magnitudes are so small that they are not visible on the table. These results lend credence to political economy and other arguments for uniform tariffs, for example as argued by Panagariya and Rodrik (1993)
@@
(Model without Services and with Input-Output Linkages)
@@
$\tau^{*}$ $10.35$ $10.40$ $11.31$ $10.35$ $10.39$ $13.94$
$\Delta W(\%)$ $0.087$ $0.071$ $-0.155$ $0.087$ $0.071$ $-0.155$
It is possible to exhaustively verify that our genetic algorithm converges to a Nash equilibrium when we have uniform tariffs. We do this by discretizing the grid of uniform tariffs the same as in Section A, and then evaluating welfare for each point on the grid. Figure B.1 presents the change in welfare for each tariff rate, assuming that their partner country sets their uniform tariff rate at the Nash equilibrium rate reported in Table B.2. Welfare is reported as consumption equivalent variation relative to the baseline in which both countries set tariffs at their Nash equilibrium levels. Welfare declines as tariff rates further depart from their Nash value, and this pattern continues well beyond the bounds displayed the graph. This exercise establishes that neither country has an incentive to deviate from the Nash equilibrium, and that our genetic algorithm converges to the correct solution.
@@ B.1.1 Welfare Impact of Proposed Tariff Hikes
Table B.3: Welfare Changes ($\Delta W$%) from Proposed Uniform Tariffs,
@@ B.2 Bilateral Trade Imbalances from Unobserved Trade Costs
In the main calibration in the paper, we assume that there are no unobserved iceberg trade costs, i.e. $t_{ij}^{s}=1$ in equation (17), which allows us to uniquely identify $\gamma_{i,j}^{s}$. In this section, we instead assume that unoberved iceberg trade costs are responsible for the bilateral trade imbalances, and we calibrate the model to match the data under this assumption. We then recompute the optimal tariff rates and welfare implications of a trade war between the United States and China.
@@
Rearranging equation (19), shows that prices of differientated output are given by
@@
We restrict $t_{ij}^{s}\geq 1$, which ensures it cannot be less costly to deliver one unit of output to a foreign country than to deliver it domestically. We further restrict $t_{ij}^{s}\leq 100$ to place an upper bound on iceberg trade costs. These restrictions are meant to prevent implausible trade costs, however, they can be relaxed if desired. These restrictions mean, for instance, that the model will interpret zeros (origin-destination-sector tuples with zero trade) as arising due to differences in preferences, rather than unobserved trade costs making it infinitely expensive to deliver output in a given sector to a specific destination. Eliminating these restrictions would force iceberg costs to fully account for global imbalances. Following calibrating iceberg trade costs, we are able to plug them into equation (26) and proceed through the remaining steps of the calibration in Section 5.
@@
B.3 Eliminating Aggregate Trade Imbalances
In static trade models, there are two common strategies to deal with aggregate trade imbalances. The first, as we did in Section 5 is to calibrate the model to match observed aggregate trade imbalances, and include the trade deficit, $D$, as a transfer in the consumer budget constraint, in equation (13). The other way to handle aggregate trade imbalances is to rebalance imports or exports for each country with the rest of the world, creating synthetic trade flows, $\widetilde{p_{{\scriptscriptstyle\text{ROW}},j}^{s}y_{{\scriptscriptstyle\text{ROW}},j}^{s}}$, that ensure trade is balanced for each country on aggregate. We implement this second method by scaling exports to the rest of the world proportionally to eliminate imbalances. Specifically, we construct,

### v2

@@ B.1 Uniform Tariff Rates Across Sectors
Our methodology can be adapted to evaluate the implications of tariff rate regimes that restrict tariff rates to be uniform across sectors. To do so, we equalize tariff rates across sectors by setting $\tau_{USA,CHN}^{s}=\tau_{USA,CHN}$, and apply our prior methodology to compute optimal unilateral and Nash equilibrium tariffs, given this restriction. Table B.2 reports optimal unilateral and Nash equilibrium tariffs when tariffs are restricted to be uniform for both the United States and China. The results are similar to those in Table 2. Furthermore, we reach the same conclusion that the United States gains from a trade war relative to the pre-trade war baseline, but experiences small welfare losses relative to a free-trade baseline.
@@
Table B.3 reports optimal unilateral and Nash equilibrium tariffs for the United States with and without the restriction that China is forced to enact uniform tariffs itself. These results show that — for the specification with input-output linkages, but without services — uniformity achieves a large majority of the potential welfare gains from optimal tariff setting compared to Table 5, where we allow tariffs to vary across sectors. Notably, the welfare implications are nearly identical regardless of whether we restrict China to implementing uniform tariffs or not. Although China does have higher welfare, and the United States has lower welfare, under the Nash equilibrium when China is not restricted to uniform tariffs, the magnitudes are so small that they are not visible on the table. These results lend credence to political economy and other arguments for uniform tariffs, for example as argued by Panagariya and Rodrik (1993).
@@
$\tau^{*}$ $9.10$ $9.28$ $14.38$ $9.10$ $9.31$ $14.77$
$\Delta W(\%)$ $0.038$ $-0.004$ $-0.133$ $0.038$ $-0.004$ $-0.133$
It is possible to exhaustively verify that our genetic algorithm converges to a Nash equilibrium when we have uniform tariffs. We do this by discretizing the grid of uniform tariffs the same as in Section C, and then evaluating welfare for each point on the grid. Figure B.1 presents the change in welfare for each tariff rate, assuming that their partner country sets their uniform tariff rate at the Nash equilibrium rate reported in Table B.3. Welfare is reported as consumption equivalent variation relative to the baseline in which both countries set tariffs at their Nash equilibrium levels. Welfare declines as tariff rates further depart from their Nash value, and this pattern continues well beyond the bounds displayed the graph. This exercise establishes that neither country has an incentive to deviate from the Nash equilibrium, and that our genetic algorithm converges to the correct solution.
@@ B.1.1 Welfare Impact of Proposed Tariff Hikes
Table B.4: Welfare Changes ($\Delta W$%) from Proposed Uniform Tariffs
@@ B.2 Bilateral Trade Imbalances from Unobserved Trade Costs
In the main calibration in the paper, we assume that there are no unobserved iceberg trade costs, i.e. $t_{ij}^{s}=1$ in equation (18), which allows us to uniquely identify $\gamma_{i,j}^{s}$. In this section, we instead assume that unobserved iceberg trade costs are responsible for the bilateral trade imbalances, and we calibrate the model to match the data under this assumption. We then recompute the optimal tariff rates and welfare implications of a trade war between the United States and China.
@@
Rearranging equation (20), shows that prices of differentiated output are given by
@@
We restrict $t_{ij}^{s}\geq 1$, which ensures it cannot be less costly to deliver one unit of output to a foreign country than to deliver it domestically. We further restrict $t_{ij}^{s}\leq 100$ to place an upper bound on iceberg trade costs. These restrictions are meant to prevent implausible trade costs, however, they can be relaxed if desired. These restrictions mean, for instance, that the model will interpret zeros (origin-destination-sector tuples with zero trade) as arising due to differences in preferences, rather than unobserved trade costs making it infinitely expensive to deliver output in a given sector to a specific destination. Eliminating these restrictions would force iceberg costs to fully account for global imbalances. Following the calibration of iceberg trade costs, we are able to plug them into equation (28) and proceed through the remaining steps of the calibration in Section 5.
@@
B.3 Alternative Method for Eliminating Aggregate Trade Imbalances
In static trade models, there are two common strategies to deal with aggregate trade imbalances. The first, as we did in Section 5 is to calibrate the model to match observed aggregate trade imbalances, and include the trade deficit, $D$, as a transfer in the consumer budget constraint, in equation (14). The other way to handle aggregate trade imbalances is to focus on scenarios in which there are no aggregate trade imbalances. In the main paper, we accomplished this by calibrating the model to the data, and then solving for a counterfactual equilibrium with $D_{i}=0$, and using this counterfactual as our baseline. An alternative approach for eliminating aggregate deficits is to rebalance imports or exports for each country with the rest of the world, creating synthetic trade flows, $\widetilde{p_{{\scriptscriptstyle\text{ROW}},j}^{s}y_{{\scriptscriptstyle\text{ROW}},j}^{s}}$, that ensure trade is balanced for each country on aggregate. We implement this second method by scaling exports to the rest of the world proportionally to eliminate imbalances. Specifically, we construct,